OpenStudy (anonymous):

Use the law of cosines to find the value of 2*3*4 cos theta.

OpenStudy (anonymous):



OpenStudy (anonymous):

by the cosine rule: A^2 = B^2 + C^2 - 2BC cos A', where A, B, C are the sides and cos A' is the cosine of the angle which is directly opposite a specified side. In the diagram, cos A' = cos (theta), A = 2, B = 3, C = 4. Subbing in: 2^2 = 3^2 + 4^2 - 2(3)(4) cos (theta) 2(3)(4) cos (theta) = 3^2 + 4^2 - 2^2 = ?

OpenStudy (anonymous):

21?

OpenStudy (anonymous):

@dauspex

OpenStudy (anonymous):

21 IS CORRECT!!!

OpenStudy (anonymous):

Yes